I am lost with this question. solve by the substitution method, and determine wehter the equation is dependent, independent or inconsistent.
3x-6y=5
2y=4x-6

Answer by Nate(3500)

2y=4x-6 original equation
y=2x-3 get 'y' by itself
plug y in the next equation:
3x-6y=5
3x - 6(2x-3)=5
3x-12x+18=5
-9x=-13
x = 13/9
plug in 'x' for one of the equation(s):
3x-6y=5
3(13/9)-6y=5
13/3 - 6y = 15/3
(-6y=2/3)(-1/6)
y = -1/9